# Accepted for BENEVOL24, the 23rd Belgium-Netherlands Software Evolution Workshop

### v2

@@ 1 Introduction
We provide an overview of the literature on how OSS package reuse practices have evolved in recent years. We propose a catalogue of challenges in OSS package dependency networks and beyond, and present recent empirical research to understand and address each of these challenges. It can be used as a starting point for junior and senior researchers as well as practitioners that would like to learn more about research challenges and advances in dependency management of OSS package registries.
2 Starting point of the literature review
@@ 3 Dependency-related challenges
Breaking changes [3, 24, 25, 26, 27, 28, 29]
@@
Insecure and vulnerable dependencies [8, 17, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]
@@ 3.1 Outdated dependencies
3.2 Breaking changes
@@
Keeping dependencies up to date requires a lot of effort from developers, because of the rapid pace at which packages are updated, but also because these updates can lead to incompatible dependencies due to the introduction of breaking changes. Bogart et al. [3] investigated the policies and practices of making and facing breaking changes in 18 software ecosystems. They observed that maintainers are frequently exposed to breaking changes, and that ecosystems differ in their approaches to breaking changes. Venturini et al. [29] analysed dependency updates in 384 npm packages and found that 11.7% of them are impacted by breaking changes when upgrading dependencies to newer versions. They also found that 44% of the dependency upgrades leading to a breaking change were supposed to be backward compatible and investigated the time it takes to address these backward incompatible changes. They observed that more than half of the backward incompatible updates are due to transitive dependencies, and that the usual mitigation strategy is either for the provider package to release a patch fixing the backward incompatibilities (usually within a week) or for the dependent packages to incorporate these backward incompatible changes in newer version of their packages (taking 4 months on average). Brito et al. [27] carried out a mixed methods empirical study to understand why and how developers introduce breaking changes in libraries. The identified reasons for breaking changes were to support new features, simplify existing APIs, and improve maintainability. They also identified a contrast between library producers and consumers in the perceived effort to overcome breaking changes. On the one side, according to the developers, the effort to adopt these breaking changes is generally limited. On the other side, nearly half of the questions related to breaking changes on StackOverflow are about how to integrate and overcome these breaking changes.
@@
Tool support. Several language-specific tools exist for helping developers detect breaking changes before an update is released to the clients. Examples include PyCompat [73], DepOwl [74] and AexPy [75] for Python, APIDiff [76], Clirr and Revapi for Java [77], and NoRegrets for JS [24, 25]. These tools analyse whether the changes made to the types used in the public API may break clients. Empirical evidence revealed that such tools are able to catch most breaking changes in practice [78]. Approaches based on type regression testing [24, 26] run the test suites of a library’s clients to detect breaking changes. While effective, it can require a lot of storage capacity and execution time. Møller and Torp [25] propose an improved variant of type regression testing for JS libraries, by automatically generating tests from a reusable API model. This approach is shown to run faster and find breaking changes in more libraries. As a complementary approach to using tests for detecting breaking changes, Zhang et al. [28] proposed and empirically validated SemBid, a tool to statically detect semantic breaking changes in third-party libraries used by Java projects by measuring semantic differences.
@@ 3.3 Versioning policies and update strategies
Proper versioning policies and update strategies, such as semantic versioning [2, 30, 35] can help in mitigating breaking changes. Semantic versioning provides an implicit convention between the library consumers (who specify the range of allowed versions) and library producers (who avoid introducing breaking changes in non-major versions). Decan and Mens [32] investigated semantic versioning compliance in Cargo, npm, Packagist and RubyGems, observing that most packages are compliant with semantic versioning. In a follow-up work [33], they found that semantic versioning is misused in packages not having reached the 1.0.0 version barrier. Ochoa et al. [5] analysed 120K library upgrades on Maven. They found that a large majority of these upgrades comply with semantic versioning and that only a minority of library consumers are affected by breaking changes.
@@ 3.4 Dependency solving
Researchers have therefore proposed better solutions. Pinckney et al. [39] proposed and evaluated Maxnpm and Pacsolve to overcome the shortcomings of the npm dependency solver. It enables customizable constraints and optimization goals, empowering developers to combine multiple objectives when installing dependencies. Other researchers have proposed more generic formally founded solutions based on constraint solving and optimisation [36]. Functional package managers, such as GNU Guix  [^1] https://guix.gnu.org and Nix  [^2] https://nixos.org, provide yet another solution by enabling the creation of separate namespaces on-the-fly, allowing to deploy incompatible packages side-by-side, thereby alleviating the need for correct and complete dependency solving.
@@ 3.6 Insecure and vulnerable dependencies
Tool support. Many tools are available to help developers detect and resolve security weaknesses in vulnerable dependencies. One of them is Dependabot, now fully integrated into GitHub. The tool issues pull requests to automatically update vulnerable dependencies and provides an effective platform for increasing awareness of dependency vulnerabilities and mitigating vulnerability threats. Alfadel et al. [51] studied the degree to which developers adopt Dependabot by investigating 2,904 OSS JS projects. They observed that a majority of security-related pull requests are accepted, often merged within a day. The severity of the dependency vulnerability and the potential risk of breaking changes were not strongly correlated with the time to merge these pull requests. Mohayeji et al. [52] empirically studied receptivity to Dependabot security updates in JS projects, observing that developers tend to delegate the task of fixing vulnerable dependencies and merge the majority of recommended security updates within several days. This is considerably faster than fixing vulnerabilities manually, which often takes up to several months. Similar to earlier research, Alfadel et al. [53] empirically analysed vulnerable dependencies in 6,546 JS applications, observing that 4.63% of them were exposed to dependencies with publicly known vulnerabilities, even if a fix was available in 90.8% of the cases. They proposed DepReveal, a tool to help developers better understand vulnerabilities in their application dependencies and to plan their project maintenance. Wang et al. [54] studied 356,283 active npm packages, observing that, in their latest release, 20% of them still introduce vulnerabilities via transitive dependencies despite the involved vulnerable packages already fixed the vulnerability for over a year. They empirically studied and distilled the remediation strategies to mitigate the fix propagation lag. Based on this, they developed Plumber, a tool to derive customised remediation suggestions for pivotal packages. The tool received positive feedback from many well-known projects.
@@
Software supply chain security. Package dependency networks can be seen as part of the wider concept of software supply chains [81, 48]. Assuring their security is crucial to avoid supply chain attacks such as the malicious update of the SolarWinds Orion monitoring software, shipping a delayed-activation trojan horse that affected thousands of organisations, including the US government. Ohm et al. [50] analysed 174 malicious software packages in npm, PyPI, and RubyGems that were used in real-world supply chain attacks. To mitigate such attacks, so-called software bills of materials (SBOM) are imposed or recommended by a US Executive Order and the EU Cyber Resilience Act to facilitate the transparent management of software dependencies, vulnerabilities, licenses, vendors, and the software supply chain. Nocera et al. [82] observed a low adoption of SBOM in public GitHub repositories of OSS projects, but with an increasing trend. Although there are significant efforts from academia and industry to facilitate SBOM development, it is still unclear how practitioners perceive SBOMs and what are the challenges of adopting SBOMs in practice. Xia et al. [83] conducted survey and interview with SBOM practitioners to understand the current state of SBOM practice, tooling support and concerns for SBOM. They identified several open challenges that need to be further studied, mitigated and addressed. In a similar vein, Stalnaker et al. [84] surveyed 138 practitioners and identified 12 major challenges concerning the creation and use of SBOMs. They propose and discuss four actionable solutions to these challenges, and present major avenues for future research and development. While SBOM has the potential to enhance vulnerability detection and facilitate license compliance, these benefits can only be realized if the SBOMs themselves are precise and correct. Yu et al. [85] examined four widely used SBOM generating tools for accuracy. They discovered that the SBOMs produced by these generators are inconsistent and contain dependency omissions, leading to incomplete and perhaps erroneous SBOMs. They consequently proposed several best practices for SBOM generation and also introduced a benchmark to steer the development of more robust SBOM generators.
@@ 3.9 Abandoned and unmaintained dependencies
Relying on reusable OSS libraries is too often taken for granted, ignoring the considerable effort required for their maintainers to keep them up to date and fixing reported issues in a timely manner. A project that depends on reusable libraries implicitly trusts the associated OSS community. This can be problematic if this community –which is often driven by volunteers– is not sufficiently active or responsive, is too small (e.g., libraries maintained by a single developer), or if libraries become unmaintained due to maintainer abandonment [63, 65]. Champion and Hill [86] studied the underproduction of software projects, where the supply of labour for maintaining them is too small to satisfy the demand of the project users. Analysing 21K+ packages with 461K+ bugs in the Debian distribution they proposed a method to identify underproductive software packages.
Miller et al. [87] quantitatively observed that abandonment is common among widely-used npm libraries, that hundreds of thousands of downstream projects were directly exposed to this abandonment, and that most of these exposed projects never remove or replace an abandoned dependency. Factors that make it more likely ro remove abandoned dependencies are more mature project government and dependency management practices, as well as the explicit public announcement of abandoned packages.
Abandoned, and therefore unmaintained, libraries can be a source of highly impactful security threats. An incident in March 2024 was the compromised XZ-Utils software compression package, impacting many Linux distributions. A malicious attacker took over maintenance of the original well-intentioned maintainer who was no longer able to fully maintain the package. A similar situation happened in November 2018 for the event-stream npm package, whose maintenance was unknowingly handed over to a malicious developer who subsequently modified the package to include code for stealing crypto-coins [88]. To avoid such incidents, it is of crucial importance to maintain a healthy and sustainable community that is able to attract and retain motivated contributors [89], and to ensure that they have the necessary financial and computing resources to maintain their code. In case packages get abandoned, the ecosystem should rely on community package maintenance organizations (CPMOs), consisting of volunteers that steward and maintain abandoned packages [66].
@@ 3.10 Incompatible licenses
Another important challenge stems from the incompatibility of OSS licenses that determine the terms and conditions to use/modify reusable libraries within one’s own software. A plethora of licenses exist (The SPDX specification supports 600+ licenses), and the use of incompatible licenses can lead to legal disputes. It is therefore essential to ensure license compatibility when reusing OSS packages, and it is a frequent reason for migrating to an alternative library [56].
Xu et al. [69] conducted a large empirical study of license incompatibilities and their remediation practices in the PyPI ecosystem for Python. They found that 7% of the package releases have license incompatibilities and 61% of them are due to transitive dependencies. They also identified 5 remediation strategies, including migrating to another library, removing the dependency, pinning versions or changing the license of the dependent package. Inspired by their findings, they proposed Silence, an approach to recommend license incompatibility remediations with minimal costs. Xu et al. [67] proposed LiDetector, a learning-based tool to detect license incompatibilities. An empirical evaluation of 1,846 projects revealed that $>72\%$ of the projects suffer from license incompatibility, including popular ones such as the MIT License and the Apache License. In a follow-up work [68], they presented LiResolver, a tool to resolve license incompatibility issues for OSS.

### v3

@@ 1 Introduction
We provide an overview of the literature on how OSS package reuse practices have evolved in recent years. We propose a catalogue of challenges in OSS package dependency networks and beyond, and present recent empirical research to understand and address each of these challenges. It can be used as a basis for junior and senior researchers as well as practitioners that would like to get a kick-start in the state-of-the-art research challenges and advances in package dependency management.
2 Starting point of the overview
@@ 3 Dependency-related challenges
Breaking changes and backward incompatibilities [3, 24, 25, 26, 27, 28, 29]
@@
Vulnerable dependencies [8, 17, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]
Supply chain attacks [53, 54, 55, 56, 57, 58, 59, 60]
@@ 3.1 Outdated dependencies
3.2 Breaking changes and backward incompatibilities
@@
Keeping dependencies up to date requires much effort from developers, because of the rapid pace of package updates, but also because these updates can lead to backward incompatibilities due to the introduction of breaking changes. Bogart et al. [3] investigated the policies and practices of making and facing breaking changes in 18 software ecosystems. They observed that maintainers are frequently exposed to breaking changes, and that ecosystems differ in their approaches to breaking changes. Through a mixed methods empirical study Brito et al. [27] analysed why and how developers introduce breaking changes in libraries. The identified reasons were to support new features, simplify existing APIs, and improve maintainability. They also identified a contrast between library producers and consumers in the perceived effort to overcome breaking changes. On the one side, according to the developers, the effort to adopt these breaking changes is generally limited. On the other side, nearly half of the questions related to breaking changes on StackOverflow are about how to integrate and overcome these breaking changes.
Venturini et al. [29] studied backward incompatibilities introduced by breaking changes when upgrading npm dependencies to newer versions. By analysing dependency updates in 384 npm packages they found that 11.7% of them lead to breaking changes, even though 44% of these dependency upgrades were meant to be backward compatible. They observed that more than half of the backward incompatible updates are due to transitive dependencies, and that the usual mitigation strategy is either for the provider package to release a patch fixing the backward incompatibilities (usually within a week) or for the dependent packages to incorporate these backward incompatible changes in newer version of their packages (taking 4 months on average). In a similar vein, Jayasuriya et al. [80] studied the Maven ecosystem by analysing 142K+ direct dependencies of 18K+ Maven artifacts. 71.6% of the dependencies were outdated, and 11.6% of the dependency upgrades applied to them resulted in breaking changes. Changes in transitive dependencies were a major factor for these breaking changes.
Recently, researchers have started to focus on so-called semantic or behavioural breaking changes. Jayasuriya et al. [81] conducted an empirical analysis on 30,548 dependencies of 8,086 Maven artifacts to identify the impact of dependency upgrades on behavioral breaking changes in the test suites of client Java projects. Only 2.30% of the dependency upgrades caused client tests to break. Zhang et al. [28] proposed and empirically validated a tool to statically detect semantic breaking changes in third-party libraries used by Java projects by measuring semantic differences.
@@
Tool support. Several language-specific tools help developers to detect breaking changes before an update is released to the clients. Examples include PyCompat [82], DepOwl [83] and AexPy [84] for Python, APIDiff [85], Clirr and Revapi for Java [86], and NoRegrets for JS [24, 25]. These tools analyse whether the changes made to the types used in the public API may break clients. Empirical evidence revealed that such tools are able to catch most breaking changes in practice [87]. Approaches based on type regression testing [24, 26] run the test suites of a library’s clients to detect breaking changes. While effective, it can require a lot of storage capacity and execution time. Møller and Torp [25] propose an improved variant of type regression testing for JS libraries, by automatically generating tests from a reusable API model. This approach is shown to run faster and find breaking changes in more libraries.
@@ 3.3 Versioning policies and update strategies
Proper versioning policies and update strategies, such as semantic versioning [2, 30, 35] can help in mitigating breaking changes. Semantic versioning provides an implicit convention between the library consumers (who specify the range of allowed versions) and library producers (who avoid introducing breaking changes in non-major versions). Decan and Mens [32] investigated semantic versioning compliance in Cargo, npm, Packagist and RubyGems, observing that most packages are compliant with semantic versioning. In a follow-up work [33], they found that semantic versioning is misused in packages not having reached the 1.0.0 version barrier. By analysing 120K library upgrades on Maven, Ochoa et al. [5] found that only a minority of library consumers are affected by breaking changes, and a large majority of the upgrades comply with semantic versioning. A study by Jayasuriya et al. [80] on dependency upgrades of outdated Maven libraries leading to breaking changes revealed that almost half of these breaking changes coincided with violations of the semantic versioning scheme.
@@ 3.4 Dependency solving
Pinckney et al. [39] proposed and evaluated Maxnpm and Pacsolve to overcome the shortcomings of the npm dependency solver. It enables customizable constraints and optimization goals, empowering developers to combine multiple objectives when installing dependencies. Other researchers have proposed more generic formally founded solutions based on constraint solving and optimisation [36].
Functional package managers, such as GNU Guix [^1] https://guix.gnu.org and Nix [^2] https://nixos.org, provide a robust solution by enforcing a declarative approach to dependency management, requiring all dependencies to be declared upfront to build or run software. This ensures that the software environment is fully defined, with each dependency explicitly specified. Additionally, these managers enable the creation of separate namespaces on-the-fly, allowing multiple versions of the same package to be installed side-by-side without any risk of incompatibility or inconsistencies. This represents a significant improvement over mainstream package managers, which often struggle to achieve similar functionality without complex workarounds or dependency resolution issues.
@@ 3.6 Vulnerable dependencies
Tool support. Many tools are available to help developers detect and resolve security weaknesses in vulnerable dependencies. One of them is GitHub’s Dependabot, which issues pull requests to automatically update vulnerable dependencies, providing an effective platform for increasing awareness of dependency vulnerabilities and mitigating vulnerability threats. Alfadel et al. [49] studied the degree to which developers adopt Dependabot by investigating 2,904 OSS JS projects. They observed that a majority of security-related pull requests are accepted, often merged within a day. The severity of the dependency vulnerability and the potential risk of breaking changes were not strongly correlated with the time to merge these pull requests. Mohayeji et al. [50] empirically studied receptivity to Dependabot security updates in JS projects, observing that developers tend to delegate the task of fixing vulnerable dependencies and merge the majority of recommended security updates within several days. This is considerably faster than fixing vulnerabilities manually, which often takes up to several months. Similar to earlier research, Alfadel et al. [51] empirically analysed vulnerable dependencies in 6,546 JS applications, observing that 4.63% of them were exposed to dependencies with publicly known vulnerabilities, even if a fix was available in 90.8% of the cases. They proposed DepReveal, a tool to help developers better understand vulnerabilities in their application dependencies and to plan their project maintenance. Wang et al. [52] studied 356,283 active npm packages, observing that, in their latest release, 20% of them still introduce vulnerabilities via transitive dependencies despite the involved vulnerable packages already fixed the vulnerability for over a year. They empirically studied and distilled the remediation strategies to mitigate the fix propagation lag. Based on this, they developed Plumber, a tool to derive customised remediation suggestions for pivotal packages. The tool received positive feedback from many well-known projects.
So-called Software Composition Analysis (SCA) tools are being increasingly adopted by practitioners to keep track of potential security risks due to vulnerable dependencies on third-party libraries. Most SCA tools are based on static and/or dynamic analysis of the project dependency graph. Imtiaz et al. [90] compared nine industry-leading SCA tools on a case study. They observed important variations in the accuracy of vulnerability reporting, suggesting that practitioners should not rely on any single tool, and that SCA tools need to achieve higher precision by avoiding false positives. Dietrich et al. [91] focused on the inverse problem of false negatives (i.e., low recall), when SCA tools miss dependencies on vulnerable components. They found empirical evidence of this for the Maven ecosystem, where somehow obfuscated clones of vulnerable components are deployed in Maven Central, but not marked as vulnerable in vulnerability databases.
3.7 Supply chain attacks
Package dependency networks are part of the wider concept of software supply chains [55, 56, 58]. Assuring their security is crucial to avoid supply chain attacks such as the malicious update of the SolarWinds Orion monitoring software, shipping a delayed-activation trojan horse that affected thousands of organisations, including the US government [92]. Ohm et al. [54] analysed 174 malicious software packages in npm, PyPI, and RubyGems that were used in real-world supply chain attacks. Nonprofit foundations such as OWASP aim to improve software security in various ways. Related to package dependencies in the software supply chain, they reported in 2022 Dependency Chain Abuse to be one of the top 10 CI/CD security risks [^3] https://owasp.org/www-project-top-10-ci-cd-security-risks/CICD-SEC-03-Dependency-Chain-Abuse. It comes with four main attack vectors: (1) Dependency confusion aims to to trick clients into downloading a malicious package from a public repository rather than a private internal package with the same name; (2) Dependency hijacking aims to compromise clients that upgrade their dependency version to a more recent, malicious version; (3) Typosquatting aims to mislead clients into using a malicious package that has a very similar name as the intended package; (4) Brandjacking aims to mislead clients in depending on malicious packages that have the characteristics of packages of a trusted brand.
@@ 3.7 Supply chain attacks
To mitigate supply chain attacks, so-called software bills of materials (SBOM) have been proposed as complete, formally structured lists of all software components present in a software product, including their licenses, versions, security vulnerabilities, and vendors. [^4] The most common SBOM formats are CycloneDX and SPDX. SBOMs are imposed or recommended by US Executive Order 14028 [93] and the EU Cyber Resilience Act [94] to facilitate the transparent management of the software supply chain. Nocera et al. [95] observed a low adoption of SBOM in public GitHub repositories of OSS projects, but with an increasing trend. Although there are significant efforts from academia and industry to facilitate SBOM development, it is still unclear how practitioners perceive SBOMs and what are the challenges of adopting SBOMs in practice. Xia et al. [96] conducted a survey and interviews with SBOM practitioners to understand the current state of SBOM practice, tooling support and concerns for SBOM. They identified several open challenges that need to be further studied, mitigated and addressed. In a similar vein, Stalnaker et al. [97] surveyed 138 practitioners and identified 12 major challenges concerning the creation and use of SBOMs. They propose and discuss four actionable solutions to these challenges, and provide suggestions for future research and development.
Tool support. SBOMs enhances vulnerability detection and facilitates license compliance (see Section 3.11), through the use of SBOM generating tools such as Trivy, Syft, Microsoft’s sbom-tool, and GitHub’s Dependency Graph. A prerequisite is that such tools achieve full precision and correctness. Unfortunately, Yu et al. [98] observed that the SBOMs generated by these tools were inconsistent and contained dependency omissions, leading to incomplete and perhaps erroneous SBOMs. They consequently proposed best practices for SBOM generation and introduced a benchmark to steer the development of more robust SBOM generators. In a similar vein, [60] observed a high variability in vulnerability reporting by different SBOM generators. Tools such as OWASP Dependency-Track [^5] https://dependencytrack.org leverage the capabilities of SBOM to identify and reduce risk in the software supply chain.
Solutions. Because secure software supply chains can be hard to attain in practice, the Linux Foundation proposes Supply chain Levels for Software Artifacts (SLSA) 666slsa.dev as a set of guidelines for supply chain security. Higher levels come with increasing security guarantees, with SLSA L3 providing the highest degree of confidence that the generated SBOM is precise, accurate and has not been tampered with. One way to achieve it is by resorting to the solution of reproducible builds [57, 59]. They ensure that, given the same source code, build environment and build instructions, bitwise identical copies of all artifacts are created. The functional package managers mentioned in Section 3.4 are instances of this solution. By requiring all dependencies to be declared upfront, they offer a robust solution to dependency management issues, ultimately supporting more reliable software supply chains.
@@ 3.10 Abandoned and unmaintained dependencies
Proper dependency management does not suffice to consider technical factors only. Human factors are equally important. Relying on reusable OSS packages all too often ignores the considerable effort required by package maintainers to keep them up to date and fixing reported issues in a timely manner [58]. By depending on reusable packages, one implicitly trusts their associated OSS community. This can be problematic if this community –which is often driven by volunteers– is not sufficiently active or responsive, is too small (e.g., packages maintained by a single developer), or if packages become unmaintained due to maintainer abandonment [69, 71]. Champion and Hill [99] studied the underproduction of software projects, where the supply of labour for maintaining them is too small to satisfy the demand of the project users. Analysing 21K+ packages with 461K+ bugs in the Debian distribution they proposed a method to identify underproductive software packages.
Miller et al. [100] quantitatively observed that abandonment is common among widely-used npm libraries, that hundreds of thousands of downstream projects were directly exposed to this abandonment, and that most of these exposed projects never remove or replace an abandoned dependency. Factors that make it more likely to remove abandoned dependencies are more mature project government and dependency management practices, as well as the explicit public announcement of abandoned packages.
Abandoned, and therefore unmaintained, libraries can be a source of highly impactful security threats. An incident in March 2024 was the compromised XZ-Utils software compression package for Linux distributions. Its original well-intentioned maintainer who was no longer able to fully maintain the package. After gaining this maintainer’s trust during a period of two years, a malicious attacker took over its maintenance, and introduced a backdoor to authorise remote code execution on affected systems. A similar situation happened in November 2018 for the event-stream npm package, whose maintenance was unknowingly handed over to a malicious developer who subsequently modified the package to include code for stealing crypto-coins [101].
These incidents underscore the importance of rigorous vetting processes to reduce the risk of social engineering that can compromise software integrity. To avoid such incidents, it is of crucial importance to maintain a healthy and sustainable community that is able to attract and retain motivated contributors [102], and to ensure that they have the necessary financial and computing resources to maintain their code. In case packages get abandoned, the ecosystem should rely on community package maintenance organizations (CPMOs), consisting of volunteers that steward and maintain abandoned packages [72].
@@ 3.11 Incompatible licenses
A final challenge stems from the incompatibility of OSS licenses that determine the terms and conditions to use or modify reusable libraries within one’s own software. A plethora of licenses exist [^7] The SPDX open standard for SBOM supports 600+ licenses, see https://spdx.org/licenses/, making it increasingly challenging for developers to select an appropriate license for their projects and to ensure that they are complying with the terms of those licenses. Given that the use of incompatible licenses can lead to legal disputes, it is essential to ensure license compatibility when reusing OSS packages, and it is a frequent reason for migrating to an alternative library [62].
Xu et al. [75] conducted an empirical study of license incompatibilities and their remediation practices in the PyPI ecosystem for Python. They found that 7% of the package releases have license incompatibilities and 61% of them are due to transitive dependencies. They also identified five remediation strategies, including migrating to another library, removing the dependency, pinning versions or changing the license of the dependent package. Inspired by their findings, they proposed Silence, an approach to recommend license incompatibility remediations with minimal costs. Wu et al. [76] conducted an empirical study of license usage, incompatibility and evolution in 33M+ packages across five package registries (Maven, npm, PyPI, RubyGems and Cargo), observing both similarities and differences in license usage across the five registries.
Tool support. In order to detect and resolve incompatible licenses, automated tools are needed. Xu et al. [73] proposed LiDetector, a learning-based tool to detect license incompatibilities. An empirical evaluation of 1,846 projects revealed that $>72\%$ of the projects suffer from license incompatibility, including popular ones such as the MIT License and the Apache License. In a follow-up work [74], they presented LiResolver, a tool to resolve license incompatibility issues for OSS.